# Scalable and Convergent Generalized Power Iteration Precoding for Massive MIMO Systems

### 04 Mar 2026 v1

@@ Appendix B Proof of Proposition 2
We prove Proposition 2 by analyzing the KKT conditions of the problem (30). For any nontrivial stationary point ${\bf f}_{k}$ and Lagrange multiplier $\lambda$, the following KKT conditions hold:
@@
$\frac{\partial}{\partial{\bf f}^{\sf H}_{k}}\bar{R}^{\sf lb}_{k}({\bf F})=A^{-1}_{k}({\bf F})(\hat{{\bf h}}^{\sf H}_{k}\hat{{\bf h}}_{k}{\bf f}_{k}+\boldsymbol{\Phi}_{k}{\bf f}_{k})-B^{-1}_{k}({\bf F})\boldsymbol{\Phi}_{k}{\bf f}_{k}.$ (83)
where $A_{k}({\bf F})=\sum_{i=1}^{K}|\hat{{\bf h}}^{\sf H}_{k}{\bf f}_{i}|^{2}+\sum_{i=1}^{K}{\bf f}_{i}^{\sf H}\boldsymbol{\Phi}_{k}{\bf f}_{i}+\frac{\sigma^{2}}{P}$ and $B_{k}({\bf F})=\sum_{i=1,i\neq k}^{K}|\hat{{\bf h}}^{\sf H}_{k}{\bf f}_{i}|^{2}+\sum_{i=1}^{K}{\bf f}_{i}^{\sf H}\boldsymbol{\Phi}_{k}{\bf f}_{i}+\frac{\sigma^{2}}{P}$. We then calculate the derivative of $\sum_{j=1,j\neq k}^{K}\bar{R}^{\sf lb}_{j}({\bf F})$ with respect to ${\bf f}^{\sf H}_{k}$, i.e.,
$\sum_{j=1,j\neq k}^{K}\frac{\partial}{\partial{\bf f}^{\sf H}_{k}}\bar{R}^{\sf lb}_{j}({\bf F})=$ $\sum_{j=1,j\neq k}^{K}C^{-1}_{j}({\bf F})(\hat{{\bf h}}^{\sf H}_{j}\hat{{\bf h}}_{j}{\bf f}_{k}+\boldsymbol{\Phi}_{j}{\bf f}_{k})$
$-\sum_{j=1,j\neq k}^{K}D^{-1}_{j}({\bf F})(\hat{{\bf h}}^{\sf H}_{j}\hat{{\bf h}}_{j}{\bf f}_{k}+\boldsymbol{\Phi}_{j}{\bf f}_{k}),$ (84)
where $C_{j}({\bf F})=\sum_{i=1}^{K}|\hat{{\bf h}}^{\sf H}_{j}{\bf f}_{i}|^{2}+\sum_{i=1}^{K}{\bf f}_{i}^{\sf H}\boldsymbol{\Phi}_{j}{\bf f}_{i}+\frac{\sigma^{2}}{P}$ and $D_{j}({\bf F})=\sum_{i=1,i\neq k}^{K}|\hat{{\bf h}}^{\sf H}_{j}{\bf f}_{i}|^{2}+\sum_{i=1}^{K}{\bf f}_{i}^{\sf H}\boldsymbol{\Phi}_{j}{\bf f}_{i}+\frac{\sigma^{2}}{P}$. We assume the contrary that $\lambda\!=\!0$. Thus, the following stationarity condition holds:
@@
Using $(v+c)^{-1}=c-(v+c)^{-1}$, where $v$ and $c$ are respectively scalar variable and constant value, we have
@@
Note that the equation in (B) holds if and only if $\hat{{\bf h}}^{\sf H}_{k}{\bf f}_{k}\!=\!0$ for all $k$. Therefore, we define the optimal precoding vector ${\bf f}^{\star}_{k}$ as
${\bf f}^{\star}_{k}=$ $\frac{1}{\lambda^{\star}}\bigg(A^{-1}_{k}({\bf F})\cdot\hat{{\bf h}}^{\sf H}_{k}\hat{{\bf h}}_{k}{\bf f}_{k}+\sum_{j=1,j\neq k}^{K}\tilde{A}^{-1}_{j}({\bf F})\cdot\hat{{\bf h}}^{\sf H}_{j}\hat{{\bf h}}_{j}{\bf f}_{k}$
$+\sum_{j=1}^{K}\tilde{A}^{-1}_{j}({\bf F})\cdot\boldsymbol{\Phi}_{j}{\bf f}_{k}\bigg),$ (88)
where $\tilde{A}_{j}({\bf F})=|\hat{{\bf h}}^{\sf H}_{j}{\bf f}_{j}|^{2}+\sum_{i=1}^{K}{\bf f}_{i}^{\sf H}\boldsymbol{\Phi}_{j}{\bf f}_{i}+\frac{\sigma^{2}}{P}$. It implies that any nontrivial stationary point ${\bf f}_{k}$ of problem in (30) must lie in the range space of estimated DL channel $\hat{{\bf H}}$ and all error covariance matrices for users ${{\boldsymbol{\Phi}}}=\left[{\boldsymbol{\Phi}}_{1},...,{\boldsymbol{\Phi}}_{K}\right]$, and the corresponding Lagrange multiplier $\lambda$ is positive. This completes the proof. ∎

### 12 Mar 2026 v2

@@ Appendix B Proof of Proposition 2
We prove Proposition 2 by analyzing the KKT conditions of the problem (30). Specifically, we establish that the Lagrange multiplier $\lambda$ must be strictly positive by using a proof by contradiction. For any nontrivial stationary point ${\bf f}_{k}$ and Lagrange multiplier $\lambda$, the following KKT conditions hold:
@@
$\frac{\partial}{\partial{\bf f}^{\sf H}_{k}}\bar{R}^{\sf lb}_{k}({\bf F})=A^{-1}_{k}({\bf F})(\hat{{\bf h}}_{k}\hat{{\bf h}}^{\sf H}_{k}{\bf f}_{k}+\boldsymbol{\Phi}_{k}{\bf f}_{k})-B^{-1}_{k}({\bf F})\boldsymbol{\Phi}_{k}{\bf f}_{k},$ (83)
where $A^{-1}_{k}({\bf F})=\left(\sum_{i=1}^{K}|\hat{{\bf h}}^{\sf H}_{k}{\bf f}_{i}|^{2}+\sum_{i=1}^{K}{\bf f}_{i}^{\sf H}\boldsymbol{\Phi}_{k}{\bf f}_{i}+\frac{\sigma^{2}}{P}\right)^{-1}$ and $B^{-1}_{k}({\bf F})=\left(\sum_{i=1,i\neq k}^{K}|\hat{{\bf h}}^{\sf H}_{k}{\bf f}_{i}|^{2}+\sum_{i=1}^{K}{\bf f}_{i}^{\sf H}\boldsymbol{\Phi}_{k}{\bf f}_{i}+\frac{\sigma^{2}}{P}\right)^{-1}$. Note that the constant factor $1/\ln 2$ arising from the derivative of $\log_{2}(\cdot)$ is absorbed into the Lagrange multiplier $\lambda$ for simplicity. We then calculate the derivative of $\sum_{j=1,j\neq k}^{K}\bar{R}^{\sf lb}_{j}({\bf F})$ with respect to ${\bf f}^{\sf H}_{k}$, i.e.,
$\sum_{j=1,j\neq k}^{K}\frac{\partial}{\partial{\bf f}^{\sf H}_{k}}\bar{R}^{\sf lb}_{j}({\bf F})=$ $\!\sum_{j=1,j\neq k}^{K}\!A^{-1}_{j}({\bf F})(\hat{{\bf h}}_{j}\hat{{\bf h}}^{\sf H}_{j}{\bf f}_{k}+\boldsymbol{\Phi}_{j}{\bf f}_{k})$
$-\!\sum_{j=1,j\neq k}^{K}\!B^{-1}_{j}({\bf F})(\hat{{\bf h}}_{j}\hat{{\bf h}}^{\sf H}_{j}{\bf f}_{k}+\boldsymbol{\Phi}_{j}{\bf f}_{k}).$ (84)
To proceed with the contradiction, we assume the contrary that $\lambda\!=\!0$. Consequently, the following stationarity condition holds:
@@
Using $v(v+c)^{-1}=1-c(v+c)^{-1}$, where $v$ and $c$ are a scalar variable and a constant value, respectively, we can represent the equation in (B) as
@@
Note that the equation in (B) holds if and only if $\hat{{\bf h}}^{\sf H}_{k}{\bf f}_{k}\!=\!0$ for all $k$. This condition implies zero desired signal power for all users, yielding a zero sum SE. This trivial solution contradicts the premise that ${\bf f}_{k}$ is a nontrivial stationary point. Thus, the optimal Lagrange multiplier must satisfy $\lambda^{\star}>0$. Consequently, we define the optimal precoding vector ${\bf f}^{\star}_{k}$ as
${\bf f}^{\star}_{k}=$ $\frac{1}{\lambda^{\star}}\Bigg(\sum_{j=1}^{K}A^{-1}_{j}({\bf F})\hat{{\bf h}}_{j}\hat{{\bf h}}^{\sf H}_{j}{\bf f}_{k}-\sum_{j=1,j\neq k}^{K}B^{-1}_{j}({\bf F})\hat{{\bf h}}_{j}\hat{{\bf h}}^{\sf H}_{j}{\bf f}_{k}=$
$+\sum_{j=1}^{K}\left(A^{-1}_{j}({\bf F})-B^{-1}_{j}({\bf F})\right)\boldsymbol{\Phi}_{j}{\bf f}_{k}\Bigg).$ (88)
It implies that any nontrivial stationary point ${\bf f}_{k}$ of the problem in (30) must lie in the range space of the estimated channel $\hat{{\bf H}}$ and all error covariance matrices for users ${{\boldsymbol{\Phi}}}=\left[{\boldsymbol{\Phi}}_{1},...,{\boldsymbol{\Phi}}_{K}\right]$, and the corresponding Lagrange multiplier $\lambda$ is positive. This completes the proof. ∎